Calculate the size of the force on Q1 due to Q2. Data: Q1-1.40 10-9 C; Q2-6.30 109 C.Use the scale below to measure distances! 02 Q1 cr 2 3 4 5 6 789 10 Submit Answer Tries 0/10
Answer #1

From the scale -

Horizontal distance between Q1 and Q2, x = 9.5 cm - 3.5 cm = 4.0 cm

Here, vertical axis is not shown in the enclosed picture. It must be given in your problem.

I am just considering the vertical distance between the two charges, y = 3.0 cm

So, distance between the two charges Q1 and Q2, r = sqrt[x^2 + y^2]

= sqrt[4.0^2 + 3.0^2] = 5.0 cm = 0.05 m

Now, expression for the electrostatic force by the charge Q2 on the charge Q1 is given as -

F = k*Q1*Q2 / r^2 = (9.0 x 10^9 x 1.40 x 10^-9 x (-6.30 x 10^-9)) / 0.05^2

= -3.175 x 10^-5 N

Negative sign shows that there will be force of attraction between the two charges.